At a particular time of day, the Sun's elevation angle is 57.5° above the horizon. For a diver who is under water and looking up toward the surface, what is the apparent elevation angle of the Sun?
Answer #1

Here we have given that,

Angle form the horizontal = 51 °

So that,

Angle of emergence e = 90-51 = 39°

So that

Relationship for apparent height is given as,

Refractive index of water uw = 1.33

Sine('e)= u× Sin39°

So that apparent elevation Angle will be ,

E' = 90 - e'

E' = 90 -  arcSin(1.33×Sin39°)

E'= 33.17°


answered by: ANURANJAN SARSAM